# A time warping model for seasonal data with application to age estimation from narwhal tusks

### v1

@@ 3.3 SMC algorithm
Inside Algorithm 1 we apply SMC to filter out the hidden process $\xi$. The SMC algorithm provides a set of $n_{p}$ particles $(\xi^{(j)})_{j=1,\ldots,n_{p}}$ and weights $(W^{(j)})_{j=1,\ldots,n_{p}}$ approximating the conditional smoothing distribution $p(\xi_{0},\ldots,\xi_{n}|y_{0},\ldots,y_{n};\theta)$ (Del Moral, Doucet and Jasra, 2006b; Doucet, De Freitas and Gordon, 2001). The SMC algorithm is an iterative algorithm. At each iteration in time, particles are sampled from a proposal distribution denoted $q(\xi_{i}|\xi_{i-1},y_{i},y_{i-1})$. To ease notation, we denote $\xi_{0:i}=(\xi_{0},\ldots,\xi_{i})$ and likewise for $y_{0:i}$.
We choose the exact transition density $p(\xi_{i}|\xi_{i-1})$ introduced in Section 1 as proposal $q$. This choice gives a simplified weight equal to
@@
The algorithm is presented in Algorithm 2 for a given value $\theta$. It provides an empirical measure which is an approximation to the smoothing distribution $p(\xi_{0:n}|y_{0:n},\theta)$. A draw from this empirical distribution can be obtained by sampling an index $j$ from a multinomial distribution with probabilities $(W^{1}_{n},\ldots,W^{n_{p}}_{n})$.
Data: $(X,Y)$, $\theta=({A},{B}_{0},{a},{b},{\sigma}^{2},{\beta},\rho,{\omega}^{2},{r}_{1},\ldots,{r}_{K})$
@@
Initialize $n_{p}\times n+1$ matrices $\xi$ and weights $W$, where $n_{p}$ is the number of particles; For $j=1,\ldots,n_{p}$, set $\xi^{(j)}_{0}=a$ and $W^{(j)}_{0}=1/n_{p}$;
@@
       Set $\lambda_{i}=2\rho c\,\xi^{(j)}_{i-1}$ for $j=1,\ldots n_{p}$;
       Update particles: $\xi^{(j)}_{i}\sim\frac{1}{2c}\chi^{\prime 2}_{\nu}(\lambda_{i})$ for $j=1,\ldots n_{p}$;
       Compute weights: $W^{(j)}_{i}=p(y_{j}\,|\xi^{(j)}_{i},\theta)$ for $j=1,\ldots n_{p}$ ;
       Normalize weights: $W^{(j)}_{i}=W^{(j)}_{i}/\sum_{i=1}^{n_{p}}W_{i}^{(j)}$ for $j=1,\ldots n_{p}$ ;
       Resample particles by drawing $n_{p}$ indices from the set $\{1,\ldots,n_{p}\}$ with corresponding probabilities $W_{i}^{(1)},\ldots,W_{i}^{(n_{p})}$. Denote the realizations $\{I_{1},\ldots,I_{n_{p}}\}$ and set $\xi_{0:i}^{(j)}:=\xi_{0:i}^{(I_{j})}$;
@@
Algorithm 2 Sequential Monte Carlo (SMC) for the S step
3.4 Estimation of the parameters $(r_{k})_{k=1,\ldots,K}$
The parameters $r_{k}$ are updated heuristically after filtering the hidden process, as outlined in Algorithm 3. The main idea of the algorithm is to exploit the fact that $r_{k}$ accounts for whatever is missing between the model with constant amplitudes and with varying amplitudes (equation (3)) at peak $k$:
$r_{k}=f(x_{k}^{*})-\underbrace{\left(A\sin(g(x_{k}^{*})+b)-B_{0}\cos(2g(x_{k}^{*})+2b)\right)}_{:=f_{0}(x_{k}^{*})}$ (18)
We approximate $f(x_{k}^{*})$ by considering the median of a proportion of neighbors. This involves considering a fraction $\zeta=0.05$ of observations surrounding each signal peak $k$ ($5\%$ above and $5\%$ below), restricted to the interval between the preceding and succeeding valleys. The average of these observations provides an estimate of the peak size $\hat{y}_{k}^{*}\approx f(x_{k}^{*})$. The displacement $\hat{r}_{k}$ is then estimated as the difference between $\hat{y}_{k}^{*}$ and the estimated baseline amplitude at peak $k$, $\hat{f}_{0}(x_{k}^{*})$, while also imposing the restriction $B(x)>A_{0}$. Thus $\hat{r}_{k}=\max(\hat{y}_{k}^{*}-\hat{f}_{0}(x_{k}^{*}),\hat{A}-\hat{B}_{0})$.
Data: $(X,Y)$, $\xi$, $\hat{\theta}=(\hat{A},\hat{B}_{0},\hat{a},\hat{b},\hat{\sigma}^{2},\hat{\rho},\hat{\omega}^{2})$
Result: $\{\hat{r}_{1},\ldots,\hat{r}_{K}\}$
Set $\zeta=0.05$ (the fraction of neighbors to consider) ;
Step(1)   Use $\hat{g}(x)=\int_{0}^{x}\xi_{s}\,\mathrm{d}s$ and equation (6) to derive peaks $x_{1}^{*},\ldots,x_{K}^{*}$;
for $k=1$ to $K$ do
       Step(2)   Find valleys $x_{k-}^{*}$ and $x_{k+}^{*}$ (local minima of $y$) before and after each peak $x_{k}^{*}$, by solving $\hat{g}(x_{k-}^{*})=g(x_{k}^{*})-\pi/2$ and $\hat{g}(x_{k+}^{*})=g(x_{k}^{*})+\pi/2$. If no valley exists before $x_{1}^{*}$, let $x_{1-}^{*}=x_{1}$. If no valley exists after $x_{K}^{*}$, let $x_{K+}^{*}=x_{n}$;
       Step(3) Define $x_{k}^{l}=x_{k}^{*}-\max(\Delta,\zeta\,(x_{k}^{*}-x_{k-}^{*}))$ and $x_{k}^{u}=x_{k}^{*}+\max(\Delta,\zeta\,(x_{k+}^{*}-x_{k}^{*}))$;
       Step(4) Define $I_{k}=\{i\in\{0,\ldots,n\}\,|\,x_{i}\in(x_{k}^{l},x_{k}^{u})\}$. Estimate the peak size:
$\hat{y}_{k}^{*}=\frac{1}{x_{k}^{u}-x_{k}^{l}}\sum_{i\in I_{k}}y_{i}$
Step(5) Define $\hat{f}_{0}(x)$ to be $f(x)$ from eq. (3), but with the estimated $\hat{A},\hat{b},\hat{g}$ plugged in, and where $B(x)=\hat{B}_{0}$. Our estimate becomes $\hat{r}_{k}=\max(\hat{y}_{k}^{*}-\hat{f}_{0}(x_{k}^{*}),\hat{A}-\hat{B}_{0})$.
@@ 3.4 Estimation of the parameters $(r_{k})_{k=1,\ldots,K}$
Algorithm 3 Estimation of $r_{k},\,k=1,\ldots,K$
In the next two sections we outline the initialization details of step (1) of Algorithm 1 and discuss the bootstrap procedure used to estimate the confidence intervals of the estimates. We end with a section on model validation.
3.5 Parameter initialization
Proper initialization will assist SAEM in convergence towards a global optimum. This is particularly important for our problem, since the loss function of the non-linear model (3) is expected to exhibit many local maxima due to the trigonometric functions. Obtaining good initial estimates is a non-trivial task, and typically involves heuristic approaches. Here we present our chosen initialization, denoted $\hat{\theta}^{0}$. The points (1) through (6) are listed in the order they should be performed.
(1) Peak detection. In order to identify peaks in the data $y$, we apply a locally estimated scatterplot smoothing (LOESS) filter (Cleveland, Grosse and Shyu, 2017), which performs local polynomial regression around each data point by using $10\%$ of the neighboring points. Denote the resulting smoothed data vector $y_{s}$. Then apply a rolling max filter on $y_{s}$, which at each point selects the maximum value within a window of size $(n/5)\Delta$. We use the rolling max function in the R-package Zoo (Zeileis and Grothendieck, 2005). Denote the resulting data vector $y_{max}$. Our initial guess at the peaks are
$\{{x_{1}^{*}}^{0},\ldots,{x_{L}^{*}}^{0}\}=\{x_{i}+(n/10)\Delta\,|\,y_{max}(x_{i})>y_{s}(x_{i}),i=1,\ldots,n\}.$ (19)
Note that the initial number of peaks $L$ might be different from the true number $K$.
(2) Amplitudes: $A,B_{0}$. Perform 2-means clustering (using the R function kmeans ) on the peak values $y_{s}({x_{1}^{*}}^{0}),\ldots,y_{s}({x_{L}^{*}}^{0})$ to identify two centroids, $HT$ and $LT$, representing the mean heights of the high and low peaks. Assuming $B(x)=B_{0}$, estimate $\hat{A}^{0}=\frac{HT-LT}{2}$ and $\hat{B}_{0}^{0}=\frac{HT+LT}{2}$, where high peaks are $A+B(x)$ and low peaks are $B(x)-A$.
(3) Frequency: $a$. Assume that the difference between consecutive peak positions correspond to half a cycle, then the cycle period is $T_{l}=2\times({x_{l}^{*}}^{0}-{x_{l-1}^{*}}^{0})$ for $l=2,..,L$. Estimates of the (angular) frequency is then $a_{l}=2\pi/T_{l}$ and we choose $\hat{a}^{0}=\textrm{median}(a_{2},\ldots,a_{L})$. We use the median to exclude extreme frequencies, for instance because we included too many peaks, or excluded intermediate peaks.
(4) Phase: $b$. Regress half of the data-vector $y(x_{i})$, $x_{i}<n\Delta/2$, on $f(x)$ in (3) with only $b$ as a free parameter, fixing $A=\hat{A}^{0}$, $B_{0}=\hat{B}_{0}^{0}$ and $\xi_{x}=\hat{a}^{0}$ (no stochastic process) for all $x$. We use non-linear least squares in the R-package minpack.lm (Elzhov et al., 2023). Limiting the analysis to the first half of the observations makes the assumption of a constant frequency $\xi_{x}=a$ for all $x$ less contentious. The initial estimate $\hat{b}^{0}$ is then the estimated parameter from the regression.
(5) Measurement error: $\sigma^{2}$. The variance of the measurement noise is approximated from the smoothed signal $y_{s}$ from the peak detection, setting $(\hat{\sigma}^{2})^{0}=\frac{1}{n}\sum_{i=1}^{n}(y(x_{i})-y_{s}(x_{i}))^{2}$.
(6) Hidden process parameters: $\rho,\omega$. Make a matrix of plausible values for $\rho$ and $\omega^{2}$. Specifically, we consider all 28 combinations of $\rho\in(0.35,0.45,0.55,0.65,0.75,0.85,0.95)$ and $\omega^{2}\in(0.4,0.8,1.2,1.6)\times\hat{a}^{0}\beta$, where $\beta=-\frac{1}{\Delta}\log\rho$. We then run SMC (Algorithm 2) with a fixed number of particles $n_{p}$ for each combination, giving a total of $28\,n_{p}$ SMC particles. We resample among the $28\,n_{p}$ particles at each step and select a final candidate based on the final weights. The selected particle is then our initial estimate $(\hat{\rho}^{0},(\hat{\omega}^{2})^{0})$.
@@ 3.6 Bootstrapping
Standard errors and confidence intervals for all parameters estimated using SAEM (Algorithm 1) are computed using parametric bootstrapping. The general procedure is described in Algorithm 4.
Data: Estimates $\hat{\theta}=(\hat{A},\hat{B}_{0},\hat{a},\hat{b},\hat{\sigma}^{2},\hat{\beta},\hat{\rho},\hat{\omega}^{2},\hat{r}_{1},\ldots,\hat{r}_{K})$
Result: Vector of estimates $(\hat{\theta}_{j})_{j=1,\ldots,M}$
$j\leftarrow 1$ ;
Set number of bootstrap replications $M\leftarrow M_{0}$ ;
while j ¡ M do
       Step(1)  Simulate data $y^{(j)}=(y_{1}^{(j)},\ldots,y_{n}^{(j)})$ using $\hat{\theta}$ and (1), (3), (4) and (7) ;
       [See section 4 for simulation details] ;
       Step(2)   Estimate $\hat{\theta}_{j}$ using Algorithm 1 on $y^{(j)}$ ;
       $j\leftarrow j+1$
@@
Algorithm 4 Bootstrapping standard errors and confidence intervals
In Step (1) of Algorithm 4 we sample data directly from the model (1), (3), (4) with estimated parameters $\hat{\theta}$. This means that $y_{i}^{(j)}=\hat{f}(x_{i})+\epsilon_{i}^{(j)}$ where $\epsilon_{i}^{(j)}\sim\mathcal{N}(0,\hat{\sigma}^{2})$.
@@ 3.7 Model validation
To assess the quality of our model, we consider the raw residuals defined as
$\textrm{res}_{i}=y_{i}-\hat{y}_{i}$ (20)
where $\hat{y}_{i}$ is the fitted value of observation $i$. We use the (raw) residuals to validate the mean structure and homoscedasticity assumption using the residual plot. Additionally, we compare the (raw) empirical residuals, equation (20), with simulated residuals (from data generated under the fitted model). This allows us to assess how well the model’s residuals reflect the variability and distributional assumptions implied by the model.
To address the estimator variability and bias, we computed the (vector of) relative difference(s) $(\tt{true\,parameter}-\tt{estimated\,parameter})/\tt{true\,parameter}$ for all parameters, except for the phase $b$. For the phase parameter $b$, we instead computed the smallest absolute difference on the unit circle. For the case-study, we applied Algorithm 4 and produced several bootstrap estimates $\tilde{\theta}_{j}$, $j=1,\ldots,M$. We then considered differences between the estimated and bootstrapped versions: $(\hat{\theta}-\tilde{\theta}_{j})/\hat{\theta}$. In the simulation study, where parameters $\theta_{m}$ are known a priori for each signal $m$, we considered the difference between estimated and true parameters: $(\theta_{m}-\hat{\theta}_{j})/\theta_{m}$.
@@ 4 Simulation study
A total of 400 signals was simulated, with a fixed step size equal to $\Delta x=x_{i}-x_{i-1}=0.1$. Below, we describe how parameters were selected. For now, assume all parameters are known. To obtain the hidden stochastic process, the solution to equation (4), we used the exact transition density (described in Section 2) to simulate a trajectory starting with $\xi_{0}=0$.
We then obtained $g(x_{i})=\sum_{j=1}^{i}\xi_{j}\Delta$ from which we retrieved the set (6). Finally, we sampled $K$ values for the displacement parameters $r_{k}$ assuming a truncated normal distribution $r_{k}\sim\mathcal{N}_{(A-B_{0},\infty)}(0,0.5^{2})$, which ensured that $B(x)=B_{0}+R(x)$ stayed positive and no crossing of upper and lower envelope of the signal. We obtained $R(x)$ using equation (7), and from that also $B(x)$. Finally, we sample a signal using (1) and (3).
For the first $M=200$ simulations, the parameters were fixed at values shared by all signals. These parameters were chosen to best reproduce the real signals of the case study. For each simulated signal $m$, we computed the relative differences between any true parameter $\theta_{m}$ and its estimate $\hat{\theta}_{m}$ (Figure 2) defined as $(\theta_{m}-\hat{\theta}_{m})/\theta_{m}$. We included the number of cycles to measure how well the model predicts the number of peaks in the signal. The violinplots in Figure 2 show how the estimates deviate from the true parameters relative to their scale. From the figure, we see that $\rho$ displays some variability, but without a prominent bias, suggesting that the rate of adjustment of the growth-time process could be hard to identify. Similarly, the amplitude parameters $A$, $B_{0}$ and the phase $b$ show a sizeable variability, without a clear bias. For the remaining parameters, there are no substantial bias, and a more dense concentration of the estimates around their true values. Most importantly, the baseline frequency $a$ and the number of cycles (used as a proxy for how well the model describe the periodic signal) display a low variance and a low bias.
Figure 2: Estimator variability. $M=200$ simulations with fixed parameters $n=500$, $\rho=0.9$, $\sigma^{2}=0.09$, $A=0.5$, $B_{0}=0.8$, $a=0.5$, $b=\frac{\pi}{2}$, $\omega^{2}=1$. The relative difference $(\theta_{m}-\hat{\theta}_{m})/\theta_{m},\,m=1,\ldots,200$ of parameter estimates $\hat{\theta}_{m}$ obtained from signals simulated with parameter $\theta_{m}$. Included are all parameters (except the $r_{k}$’s), and the number of cycles are computed with formula (5). For $b$ we computed the smallest absolute relative difference on the unit circle, since there are no principal direction.
For the other $M=200$ simulations, we allowed the parameters to vary between signals, in order to more deeply explore how well our method identifies the number of cycles, and to which extent it deviates over a variety of signals and parameter configurations (Figure 3). The signal length $n$ was sampled with uniform probability between one of four options: $200$ (short), $400$ (medium short), $600$ (medium long), $800$ (long). We then set the minimum number of cycles $C_{min}=2$ and the maximum number of cycles $C_{max}=\max(6,n/100\times 1.5)$, i.e., 6, 9 or 12 cycles. We let $a_{min}=2\pi C_{min}/N\Delta$ and $a_{max}=2\pi C_{max}/N\Delta$. The sampling distributions of the other parameters were uniformly distributed $\mathcal{U}$ as
$\rho\sim\mathcal{U}(0.55,0.95),\quad\beta=-\frac{1}{\Delta}\log(\rho),\quad\sigma\sim\mathcal{U}(0.1,0.5),$ (21)
$A\sim\mathcal{U}(0.4,0.6),\quad B_{0}\sim\mathcal{U}(0.6,1),\quad a\sim\mathcal{U}(a_{min},a_{max})$ (22)
$b\sim\mathcal{U}(0,2\pi),\quad\omega\sim\mathcal{U}(0.3,0.8)\sqrt{2a\beta}$ (23)
In this part of the simulation, the standard deviation of the displacement parameters $r_{k}$ was sampled from the uniform distribution $\mathcal{U}(0.3,0.7)$. The support of the uniform distributions were inspired by estimates from the case-study signals, and by considering the visual appearance of different parameter configurations. The result of this second part of the simulation is displayed in Figure 3 which shows that $92\%$ of the estimates fall within 1 cycle from the true number of cycles, and less than $1\%$ deviate by more than 3 cycles.
Figure 3: Estimated number of cycles against true number of cycles. Red, blue and green ribbons enclose estimates of number of cycles from simulated data sets with less than 1, 2, and 3 cycles of deviation from the true number of cycles, respectively.
@@
An example of a simulated signal is presented in the top plot of Figure 4, which also displays the model fit. In the second row of Figure 4 we have provided visualizations of the hidden process $(\xi_{x})_{x\geq 0}$ associated with the simulated signal, both the actual process and the estimated version. The third row of Figure 4 displays the growth-time process $\hat{g}(x)$, along with its estimated counterpart, obtained through the integration of $(\hat{\xi}_{x})_{x\geq 0}$. The last row shows the amplitude processes $R(x)$.
@@
Figure 4: Simulated and fitted signal. Simulated signals are in black, fitted signals in blue. The data set was simulated using $n=500$, $\Delta=1$, $\rho=0.82$, $\beta=0.2$, $\sigma^{2}=0.04$, $A=0.5$, $B_{0}=0.8$, $a=0.05$, $b=\pi/2$, $\omega^{2}=0.01$. The amplitude displacement parameters $r_{k}$ were simulated from a zero-mean normal distribution with variance 0.25. The SAEM-SMC algorithm used $n_{p}=500$ number of particles. A. Simulated signal and fitted signal. B. Simulated hidden process $(\xi_{x})_{x\geq 0}$ and estimate $(\hat{\xi}_{x})_{x\geq 0}$. C. Simulated growth time process $g(x)$ and estimate $\hat{g}(x)$. D. Amplitude correction function $R(x)$ and estimate $\hat{R}(x)$. Gray dashed lines are (true) signal peaks. Blue dashed lines are estimated signal peaks.

### v2

@@ 3.5 SMC algorithm
In Algorithm 1 we apply SMC to filter out the hidden process $\xi$. The SMC algorithm provides a set of $n_{p}$ particles $(\xi^{(j)})_{j=1,\ldots,n_{p}}$ and weights $(W^{(j)})_{j=1,\ldots,n_{p}}$ approximating the conditional smoothing distribution $p(\xi_{0},\ldots,\xi_{n}|y_{0},\ldots,y_{n};\theta)$ (Del Moral, Doucet and Jasra, 2006b; Doucet, De Freitas and Gordon, 2001). The SMC algorithm is an iterative algorithm. At each iteration, particles are sampled from a proposal distribution $q(\xi_{i}|\xi_{i-1},y_{i},y_{i-1})$. To ease notation, we denote $\xi_{0:i}=(\xi_{0},\ldots,\xi_{i})$ and likewise for $y_{0:i}$.
We choose the transition density $p(\xi_{i}|\xi_{i-1})$ (see Section 2) as proposal $q$. This choice simplifies the weight to
@@
The basic SMC algorithm is presented as pseudocode in Algorithm 2 for a given value $\theta$. It provides an empirical measure which is an approximation of the smoothing distribution $p(\xi_{0:n}|y_{0:n},\theta)$. A sample from this empirical distribution is obtained by sampling an index $j$ from a multinomial distribution with probabilities $(W^{1}_{n},\ldots,W^{n_{p}}_{n})$.
Phase offset $b$ and the growth process $g(x)$ are intrinsically connected and need to be handled jointly. We therefore propose to use SMC with a grid search adaptation, denoted SMC+, described in Algorithm 3. SMC+ uses equally spaced candidates of $b$, centered around the current estimate of $b$, and runs the basic SMC with each candidate. We consider $G=20$ equidistant values. We set a range $2w$ around the current estimate to search. Initially, we set $w=\pi$, which decreases as the Stochastic Approximation sequence in Eq. (15) decreases, resulting in a grid that collapses around the current estimate of $b$.
Data: $(X,Y)$, $\theta=({A},{B},{a},{b},{\sigma}^{2},{\beta},\rho,{\omega}^{2})$
@@
(1)  Initialize $n_{p}\times n+1$ matrices $\xi$ and weights $W$, where $n_{p}$ is the number of particles;
(2)   For $j=1,\ldots,n_{p}$: set $\xi^{(j)}_{0}=a$ and $W^{(j)}_{0}=1/n_{p}$;
@@
 (3)  Set $\lambda_{i}=2\rho c\,\xi^{(j)}_{i-1}$ for $j=1,\ldots n_{p}$;
 (4)  Update particles: $\xi^{(j)}_{i}\sim\frac{1}{2c}\chi^{\prime 2}_{\nu}(\lambda_{i})$ for $j=1,\ldots n_{p}$;
 (5)  Compute weights: $W^{(j)}_{i}=p(y_{j}\,|\xi^{(j)}_{i},\theta)$ for $j=1,\ldots n_{p}$;
 (6)  Normalize weights: $W^{(j)}_{i}=W^{(j)}_{i}/\sum_{i=1}^{n_{p}}W_{i}^{(j)}$ for $j=1,\ldots n_{p}$;
 (7)  Resample particles by drawing $n_{p}$ indices from the set $\{1,\ldots,n_{p}\}$ with probabilities $W_{i}^{(1)},\ldots,W_{i}^{(n_{p})}$. Denote the realizations $\{I_{1},\ldots,I_{n_{p}}\}$ and set $\xi_{0:i}^{(j)}=\xi_{0:i}^{(I_{j})}$;
@@
Algorithm 2 Sequential Monte Carlo (SMC)
Data: $(X,Y)$, $\theta=({A},{B},b,{a},{\sigma}^{2},{\beta},\rho,{\omega}^{2})$
Result: $\hat{b},\hat{\xi}$
(0)   Set number of candidates $G$ (grid granularity); (1)   Define half‐width
$w\;=\;\begin{cases}\pi,&m\leq m_{0},\\ \pi\times(m-m_{0})^{-0.8}&m>m_{0}.\end{cases}$
(2)   Form grid of $G$ phase zero candidates:
$b_{j}=\bigl{(}b-w+(j-1)\tfrac{2w}{G-1}\bigr{)}\bmod 2\pi,\quad j=1,\dots,G.$
for $j=1$ to $G$ do
 (3) Run SMC (Algorithm 2) with $\theta$ and $b_{j}$ replacing $b$, to obtain $\hat{\xi}^{(j)}$;
 (4) Given $\theta$, with $b_{j}$ replacing $b$, and $\hat{\xi}^{(j)}$ compute the log likelihood (9) and denote it $L_{j}$;
@@
(5)   Set $(\hat{b},\hat{\xi})=(b_{j^{*}},\xi^{(j^{*})})$ where $L_{j^{*}}\geq L_{j}$ for $j=1,\ldots,G$.
Algorithm 3 SMC+: Grid Search
@@ 3.6 Bootstrapping
Standard errors and confidence intervals for all parameters estimated using SAEM (Algorithm 1) are computed using residual bootstrapping. The general procedure to obtain a set of bootstrapped estimates is described in Algorithm 4. From these relevant summary measures can be computed.
Data: Observations $y=(y_{1},\ldots,y_{n})$ and fitted values $\hat{y}=(\hat{y}_{1},\ldots,\hat{y}_{n})$
Result: Vector of estimates $(\hat{\theta}_{m})_{m=1,\ldots,M}$
$m\leftarrow 1$ ;
(1)  Set number of bootstrap replications $M$;
(2)   Compute residuals $r_{i}=y_{i}-\hat{y_{i}}$, for $i=1,\ldots,n$;
while m ¡ M do
 (3)   Let $r_{i}^{(m)}$ be a random sample from the set of residuals $\{r_{1},r_{2},\ldots,r_{n}\}$, for $i=1,\ldots,n$;
 (4)   Construct replicate signal as $y_{i}^{(m)}=\hat{y}_{i}+r_{i}^{(m)}$, for $i=1,\ldots,n$;
 (5)   Estimate $\hat{\theta}_{m}$ using Algorithm 1 on the replicate $(y_{1}^{(m)},\ldots,y_{n}^{(m)})$ ;
 $m\leftarrow m+1$
@@
Algorithm 4 Residual Bootstrapping (RB)
@@ 3.7 Model validation
To assess the quality of the model, we use the raw residuals,
$\textrm{res}_{i}=y_{i}-\hat{y}_{i},$ (21)
where $\hat{y}_{i}$ is the fitted value of observation $i$. We validate the mean structure and homoscedasticity assumption by residual plots. Additionally, we compare the empirical residuals (21) with theoretical normally distributed errors.
To address the estimator variability and bias, we compute the (vector of) relative differences between the estimated and bootstrapped versions: $(\tilde{\theta}-\hat{\theta}_{m})/\hat{\theta}$.
@@ 4 Simulation study
In this Section, we simulate a collection of diverse signals - varying in parameter configurations and lengths - and evaluate both the accuracy of the estimates and the precision with which we recover the number of cycles (5).
A total of 800 signals were simulated, with a fixed step size equal to $\Delta=x_{i}-x_{i-1}=1$ and individual parameters $\theta_{m},m=1,\ldots,800$. To obtain the hidden stochastic process, the solution to Eq. (4), we simulated trajectories using the Euler-Maruyama scheme with step size $\Delta/100$ starting at $\xi_{0}=0$. We then obtained $g(x_{i})=g(x_{i-1})+\sum_{j=1}^{100}\xi_{(i-1)j}\Delta/100$, where $\xi_{ij}$ is the simulated value at $x=i+j/100$. Finally, we sampled a signal using Eqs. (1) and (3).
The simulation details are as follows. We simulated 200 data sets for each of 4 sample sizes; short ($n=200$), medium short ($n=400$), medium long ($n=600$) and long ($n=800$). We set $C_{\min}=2$ and $C_{\max}=\max(6,n/100\times 5)$ to be the minimum and maximum of allowed number of cycles. The upper and lower frequency is then set to $a_{\min}=2\pi C_{\min}/N\Delta$ and $a_{\max}=2\pi C_{\max}/N\Delta$. The values of the parameters were sampled from uniform distributions $\mathcal{U}$ with the following supports,
$\beta\sim\mathcal{U}(0.01,3),\quad a\sim\mathcal{U}(a_{min},a_{max}),\quad\omega\sim\mathcal{U}(0.01,0.3)$ (22)
$\sigma\sim\mathcal{U}(0.2,0.6),\quad b\sim\mathcal{U}(0,2\pi),\quad A\sim\mathcal{U}(0.5,1),$ (23)
and from this, B=1AB=1-A and $\rho=\exp(-\Delta\beta)$. The supports of the uniform distributions were chosen to best reproduce the real signals of the case study. For the process parameters in Eq. (22) we did rejection-sampling, and only accepted the triple $(a,\beta,\omega)$ if they fulfilled the Feller condition $2a\beta>\omega^{2}$ to ensure that $\xi_{x}>0$ for all $x$.
@@
A simulated signal is presented in the top plot of Figure 3, which also displays the model fit. The second row shows the hidden process $(\xi_{x})_{x\geq 0}$ underlying the simulated signal, both the true process and the estimated version. The third row displays the growth process $\hat{g}(x)$, along with its estimated counterpart.
@@
Figure 3: Simulated and fitted signal. Simulated signals are in black, fitted signals in blue. The dataset was simulated using $n=500$, $\Delta=1$, $\beta=0.07$, $\sigma=0.3$, $A=0.6$, $B=0.4$, $a=0.05$, $b=\pi/20$, $\omega^{2}=0.064$. The algorithm used $n_{p}=1500$ number of particles. A. Simulated signal and fitted signal. B. Simulated hidden process $(\xi_{x})_{x\geq 0}$ and estimate $(\hat{\xi}_{x})_{x\geq 0}$. C. Simulated growth process $g(x)$ and estimate $\hat{g}(x)$.
For each simulated signal $m$, we computed the relative difference between the true parameter $\theta_{m}$ and its estimate $\hat{\theta}_{m}$ defined as $(\hat{\theta}_{m}-\theta_{m})/\theta_{m}$, including the number of cycles. Figure 4 shows that the estimates of $a$ and $A$ are close to the true values relative to their scale, whereas the other parameters have larger variance, but without a prominent bias, maybe except for $b$. Most importantly, the number of cycles, Eq. (5), displays low variance and low bias, which is one of the key questions, since it provides an age estimate. Figure 5 shows that $91\%$ of the estimates fall within 1 cycle from the true number of cycles, and less than $4\%$ deviate by more than 3 cycles, which is mainly happening for signals with many cycles, where the peaks are more tightly concentrated.
Figure 4: Estimator variability. $M=800$ simulations equally divided into 4 groups of different length ($n=200$, $n=400$, $n=600$, $n=800$) and with random parameter configurations (see Section 4). Violinplots show the relative difference $(\theta_{m}-\hat{\theta}_{m})/\theta_{m},\,m=1,\ldots,800$ of parameter estimates $\hat{\theta}_{m}$ obtained from signals simulated with parameter $\theta_{m}$. The number of cycles are computed with Eq. (5). For $b$ we computed the smallest absolute relative difference on the unit circle. This was divided by $\pi$ to obtain a relative difference.
Figure 5: Estimated number of cycles against true number of cycles. The colored ribbons enclose estimates of number of cycles from simulated data sets with less than 1, 2, and 3 cycles of deviation from the true number of cycles.
In Section 5 we fit the model to data obtained from measurements of isotopes within narwhal tusks.